A compound contains 6.0 g of carbon and 8.0 g of oxygen. The percent composition of the compound is:
1
Expert's answer
2020-02-10T03:37:07-0500

Task:

A compound contains 6.0 g of carbon and 8.0 g of oxygen. The percent composition of the compound is:


Solution:

m(carbon) = 6.0 g;

m(oxygen) = 8.0 g

total weight of the compound = m(compound) = m(carbon) + m(oxygen);

m(compound) = 6.0 g + 8.0 g = 14.0 g;

Hence,

w(carbon) = m(carbon) / m(compound) = 6.0 / 14.0 = 0.4286 or 42.86%.

w(oxygen) = m(oxygen) / m(compound) = 8.0 / 14.0 = 0.5714 or 57.14%.


Answer: The percent composition of the compound is: 42.86% carbon and 57.14% oxygen.